OpenStudy (chunkymonkay):

Which is a equation in point-slope form for the given point and slope? Point 2, -6 Slope -3/4

OpenStudy (chunkymonkay):

@bibby @KamiBug

OpenStudy (bibby):

what does a line in point slope form look like?

OpenStudy (chunkymonkay):

a sraigt line

OpenStudy (bibby):

what does the equation look like*

OpenStudy (chunkymonkay):

y=mx+b

OpenStudy (bibby):

\(y-y_1=m(x-x_1)\) where m is the slope and \((x_1, y_1)\) is the point we want the line to pass through

OpenStudy (chunkymonkay):

so would my equation look like this? y-6=3/4(x-2)

OpenStudy (bibby):

close. \(y--6=-\frac{3}{4}(x-2)\)

OpenStudy (chunkymonkay):

oh I forgot the - sight thank you can you help me with like 4 more?

OpenStudy (bibby):

maybe

OpenStudy (chunkymonkay):

theres nothing that says y--6=-3/4(x-2) in the answers

OpenStudy (bibby):

distribute the -3/4 a nd then simplify it until it's in y=mx+b form

OpenStudy (chunkymonkay):

ok thanks

OpenStudy (bibby):

let me know what you get so I can check your work

OpenStudy (chunkymonkay):

I got y-6=-3/4(x+2)

OpenStudy (bibby):

you didn't change anything

OpenStudy (chunkymonkay):

im really confused

OpenStudy (bibby):

distribute the -3/4 into the parentheses

OpenStudy (chunkymonkay):

ok hold on

OpenStudy (chunkymonkay):

y-6=(-3/4)-x+2 that dosnt seem right to me

OpenStudy (bibby):